# ConRF: Zero-shot Stylization of 3D Scenes with Conditioned Radiation Fields

### 02 Feb 2024 v1

@@ 1 Introduction
  • We propose a novel method that leverages CLIP for zero-shot 3D scene style transfer by a single condition (i.e. image or text).
@@ 2.1 Style Transfer on Image and Video
Artistic image stylization is a long-term problem for computer vision. The primary objective is to naturally blend the content features from one image with the style features from another, resulting in a unique image that seamlessly incorporates both elements. The early traditional method is to use handmade methods to simulate style [14, 15]. With the development of deep learning, [11] is the first work that uses CNN to separate and manipulate the content and style of images for style transfer. Since then, style transfer based on neural networks has gradually evolved from single-style transfer to multi-style transfer, and now to arbitrary style transfer [6, 18, 24, 25, 28, 37, 41, 48]. In terms of video style transfer, enforcing the temporal smoothness constraints defined on the optical flow can successfully transfer the style of the video [2, 16, 40, 47, 47]. Nevertheless, all stylization methods, whether for images or videos, currently rely on a predetermined reference style, imposing substantial constraints on the creative possibilities.
@@ 2.2 Style Transfer on NeRF
Recently, neural implicit representation methods, such as NeRF [27, 34, 50, 33, 51, 52, 53, 1, 31, 8, 30], have shown great potential for high-quality rendering. NeRF leverages multi-layer perceptrons (MLPs) to implicitly model continuous scenes, leading to impressive results in view synthesis when rendering scenes. The emergence of NeRF has made significant progress in the neural network representation of 3D scenes. More recently, several works [5, 19, 36, 54, 57, 26, 4, 9] combine NeRF with neural style transfer [12] to handle artistic 3D scene stylization, [21, 13] are based on the diffusion model and can use text to stylize 3D scenes. In particular, [5, 36, 54] incorporate style loss functions from 2D images to fine-tune pre-trained NeRF models. Meanwhile, [26, 19] refine pre-trained NeRF models using a 2D style transfer model based on previous work. The approach presented in [4] employs a pre-trained 2D realistic network to constrain the realistic styles of different views and style images within a 3D scene. It’s worth noting, however, that the primary focus of this method is on transferring only the color tone of the style image. SNeRF [36] addresses the memory limitations associated with whole-image training in NeRF and enhances visual quality by employing a cyclic process of stylization and NeRF training. On the other hand, ARF [54] and Ref-NPR [56] aim to transfer detailed style features through matching features between style images and scenes. However, both SNeRF, ARF, and Ref-NPR entail a time-intensive optimization process for each reference style. While [9] and [19] leverage latent coding to achieve commendable results in style migration, their methods are limited to known styles and cannot adapt to unseen ones. [5] can achieve arbitrary style transfer by implicitly injecting style information into MLP parameters. Similar to [4], it can only transfer the color tone of the style image. Most of the neural style transfer-based methods mentioned above necessitate the use of style reference images to steer the style transfer, whereas our approach primarily focuses on neural style transfer-based methods that leverage text or image for style transfer. To accomplish text style transfer, we utilize CLIP-encoded text embeddings to manipulate the scene for style transfer, leveraging only the already aligned text images in the pre-trained CLIP latent space.
@@ 3.4 Inference process
Eq. 18 can be interpreted as the individual transfer of style to every sampling point along the ray prior to volume rendering. Ultimately, we employ a decoder to reproject these stylized features into RGB space, thereby yielding stylized novel views. In addition, our local style transfer feature is optional and becomes active only once a content text prompt is entered. Our method supports both reference image style transfer and text prompt style transfer, enabling versatile multi-combination style transfers. After inputting content text prompts, our method allows for the use of various prompt combinations to transform different aspects. It supports combinations such as text-text, image-image, and text-image for style transfers.
@@ 4.3 Ablation Studies
Golbal stylization In Fig. 8, we present a comparison between our full system and its variants, where we removed specific modules individually: A) mapping module, B) style feature loss, and C) consistency loss. As illustrated, the absence of a mapping module impedes the transfer of style from any reference image. The lack of a style feature loss compromises the quality of the transferred style, even though some elements of the reference picture’s style may still be obtained. Moreover, the inclusion of a consistency loss noticeably enhances the performance of style transfer.
Local stylization We control local style transfer through text prompts. As illustrated in Fig. 9, we conduct a comprehensive evaluation of the Multi-spatial feature extraction strategy by experimenting with different sliding window sizes. Our observations indicate that utilizing a window size of 32 leads to the generation of a notably refined mask, effectively achieving our goal of fine-grained style control in localized areas.
@@ 4.4 Limitations
Our model benefits from CLIP, while also facing certain limitations imposed by it. For example, as shown in Figure 10, we were not able to migrate the blue text style well because we used CLIP directly without making any changes to it. We expect future work to successfully address them. Additionally, our local style transfer currently only works in face-forwarding scenes.
@@ 5 Conclusion
Supplementary Material
6 Local style transfer details
The local style transfer branch is activated, we can use a content text prompt and $F_{c}$ calculate the similarity $z$:
$z=cos\left\langle F_{c},F_{\text{CLIP}}\right\rangle,$ (19)
$m=z\leq t,$ (20)
$t$ is the threshold. Thuse Eq. 18 can be replace:
7 Implementation details
Our model is trained on a single RTX 3090 GPU using approximately 80,000 samples from the WiKiArt dataset [43] as style images. During the training phase, the Adam optimizer [22] is employed with a learning rate of 1e-4. We assign the loss weights for style and content as 20 and 1, respectively, while all other loss weights are maintained at 1. For incorporating CLIP [38], we utilize the pre-trained ViT-B/32 model.
Mapping network architecture
We present the architecture of the mapping network, the details please see Fig. 11.
Figure 11: Mapping network architecture. $F_{s}$ represents the style feature, which consists of the mean and standard deviation obtained through mapping the CLIP feature ($F_{CLIP}$). Meanwhile, $\hat{F_{s}}$ denotes the stylized feature.
8 User study
We conduct a user study, following previous research [26], where we compare our approach to a 3D style transfer baseline. We engage 25 participants with diverse demographic backgrounds, including different professions, age ranges, and ethnicities, to ensure a broad perspective. Each participant is shown a series of stylizations that include original scene videos, their stylized counterparts, and videos processed using both our ConRF method and the baseline technique.
Participants are tasked with two evaluations: identifying the stylized video that most accurately reflects the style of a given image, and selecting the video that exhibits the best multi-view consistency. For a comprehensive analysis, we prepare 24 unique scene-style pairings, which we then randomly distribute into 6 categories. Each participant evaluates one category, providing us with a diverse set of opinions on the effectiveness of our approach. The aggregated findings from this study are detailed in Table 2, as referenced in our report.
Methods Stylization Score Consistency Score
Ref-NPR [56] 0.473 0.766
SNeRF [36] 0.526 0.696
ARF [54] 0.450 0.795
StyleRF [26] 0.510 0.800
Our 0.693 0.843
$\text{CLIPStyler}^{*}$ [23] 0.605 0.333
[…]
$\text{{Our}}^{*}$ 0.620 0.880
Table 2: User study. We compare ConRF with the state-of-the-art 3D style transfer baseline. (***) is using a text prompt to style transfer. The best score is red, and the second score is blue.
Figure 12: Comparison with StyleRF [26] with different views on LLFF. Here, we present the flower scene style transfer results.
9 Additional Experiment
We present more qualitative results here and video results in the supplementary material.
9.1 Quantitative results
Text prompt style transfer comparison on LLFF database
As shown in Fig. 13, we additionally perform evaluations on the LLFF [33].
Figure 13: Comparison with two SOTA style transfer methods using text prompt on LLFF. Here, we present the horns and room scenes style transfer results.
Figure 14: Local style transfer. In the main text, we present the transfer results of the text-image combination, where we show the transfer results of text-text and image-image.
Local style transfer quantitative results
As outlined in the main text, our model is versatile, supporting text-image, text-text, and image-image combinations to generate diverse stylization outcomes. The text-text and image-image combinations, which demonstrate this flexibility, are exemplified in Figure 3 of the supplementary materials (Fig. 14).
More quantitative results
To demonstrate the performance of our method, as shown in the Fig. 12, Fig. 15, Fig. 16, Fig. 17, Fig. 18, and Fig. 19, here we provide more visualization results. Additionally, our video results are also provided in the supplementary material file.
Figure 15: Style transfer result on Synthetic NeRF dataset. We present the text and image style transfer results of the drums scene.
Figure 16: Style transfer result on Synthetic NeRF dataset. We present the text and image style transfer results of the hotdog scene.
Figure 17: Style transfer result on Synthetic NeRF dataset. We present the text and image style transfer results of the mic scene.
Figure 18: Style transfer result on Synthetic NeRF dataset. We present the text and image style transfer results of the chair scene.
Figure 19: Style transfer result on Synthetic NeRF dataset. We present the text and image style transfer results of the orchids scene.

### 06 Mar 2024 v2

# https://xingy038.github.io/ConRF/

@@ 1 Introduction
  • We propose a novel method that leverages CLIP for zero-shot 3D scene artistic style transfer by a single condition (i.e. image or text).
@@ 2.1 Style Transfer on Image and Video
Artistic image stylization is a long-term problem for computer vision. The primary objective is to naturally blend the content features from one image with the style features from another, resulting in a unique image that seamlessly incorporates both elements. The early traditional method is to use handmade methods to simulate style [15, 16]. With the development of deep learning, [12] is the first work that uses CNN to separate and manipulate the content and style of images for style transfer. Since then, style transfer based on neural networks has gradually evolved from single-style transfer to multi-style transfer, and now to arbitrary style transfer [7, 19, 25, 26, 29, 38, 42, 48]. In terms of video style transfer, enforcing the temporal smoothness constraints defined on the optical flow can successfully transfer the style of the video [2, 17, 41, 47]. Nevertheless, all stylization methods, whether for images or videos, currently rely on a predetermined reference style, imposing substantial constraints on the creative possibilities.
@@ 2.2 Style Transfer on NeRF
Recently, neural implicit representation methods, such as NeRF [28, 33, 50, 35, 51, 52, 53, 1, 31, 8, 32], have shown great potential for high-quality rendering. NeRF leverages multi-layer perceptrons (MLPs) to implicitly model continuous scenes, leading to impressive results in view synthesis when rendering scenes. The emergence of NeRF has made significant progress in the neural network representation of 3D scenes. More recently, several works [5, 20, 37, 54, 57, 27, 4, 10] combine NeRF with neural style transfer [13] to handle artistic 3D scene stylization, [22, 14] are based on the diffusion model and can use text to stylize 3D scenes. In particular, [5, 37, 54] incorporate style loss functions from 2D images to fine-tune pre-trained NeRF models. Meanwhile, [27, 20] refine pre-trained NeRF models using a 2D style transfer model based on previous work. The approach presented in [4] employs a pre-trained 2D realistic network to constrain the realistic styles of different views and style images within a 3D scene. It’s worth noting, however, that the primary focus of this method is on transferring only the color tone of the style image. SNeRF [37] addresses the memory limitations associated with whole-image training in NeRF and enhances visual quality by employing a cyclic process of stylization and NeRF training. On the other hand, ARF [54] and Ref-NPR [56] aim to transfer detailed style features through matching features between style images and scenes. However, both SNeRF, ARF, and Ref-NPR entail a time-intensive optimization process for each reference style. While [10] and [20] leverage latent coding to achieve commendable results in style migration, their methods are limited to known styles and cannot adapt to unseen ones. [5] can achieve arbitrary style transfer by implicitly injecting style information into MLP parameters. Similar to [4], it can only transfer the color tone of the style image. Most of the neural style transfer-based methods mentioned above necessitate the use of style reference images to steer the style transfer, whereas our approach can leverage text or image for style transfer. To accomplish text style transfer, we utilize CLIP-encoded text embeddings to manipulate the scene for style transfer, leveraging only the already aligned text images in the pre-trained CLIP latent space. Furthermore, our work only focuses on artistic style transfer and does not have the creative capabilities based on generative models.
@@ 3.4 Inference process
Eq. 18 can be interpreted as the individual transfer of style to every sampling point along the ray prior to volume rendering. Ultimately, we employ a decoder to reproject these stylized features into RGB space, thereby yielding stylized novel views. In addition, our local style transfer feature is optional and becomes active only once a content text prompt is entered. The local style transfer branch is activated, we can use a content text prompt and $F_{c}$ calculate the similarity $z$:
$z=cos\left\langle F_{c},F_{\text{CLIP}}\right\rangle,$ (19)
$m=z\leq t,$ (20)
$t$ is the threshold. Thuse Eq. 18 can be replace:
Our method supports both reference image style transfer and text prompt style transfer, enabling versatile multi-combination style transfers. After inputting content text prompts, our method allows for the use of various prompt combinations to transform different aspects. It supports combinations such as text-text, image-image, and text-image for style transfers.
@@ 4.3 Ablation Studies
Golbal stylization
In Fig. 8, we present a comparison between our full system and its variants, where we removed specific modules individually: A) mapping module, B) style feature loss, and C) consistency loss. As illustrated, the absence of a mapping module impedes the transfer of style from any reference image. The lack of a style feature loss compromises the quality of the transferred style, even though some elements of the reference picture’s style may still be obtained. Moreover, the inclusion of a consistency loss noticeably enhances the performance of style transfer.
Local stylization
We control local style transfer through text prompts. As illustrated in Fig. 9, we conduct a comprehensive evaluation of the Multi-spatial feature extraction strategy by experimenting with different sliding window sizes. Our observations indicate that utilizing a window size of 32 leads to the generation of a notably refined mask, effectively achieving our goal of fine-grained style control in localized areas.
@@ 4.4 Limitations
Our model benefits from CLIP, while also facing certain limitations imposed by it. For example, as shown in Figure 10, we were not able to migrate the blue text style well because we used CLIP directly without making any changes to it. Our work currently only focuses on artistic style transfer and does not have creative capabilities based on generative models. We expect future work to successfully address them. Additionally, our local style transfer currently only works in face-forwarding scenes.